How do I find the maximum area for a rectangle with a fixed perimeter of 120
Found 2 solutions by rfer, stanbon:
Answer by rfer(12666)

a square is max'
120/4=30
30*30=900 sq units

Answer by stanbon(57384)

): How do I find the maximum area for a rectangle with a fixed perimeter of 120
----
Area = L*W
Perimeter = 2L + 2W
------
2L+2W = 120
L + W = 60
W = 60-L
-----
Substitute into the Area formula:
A = L(60-L)
A = 60L - L^2
-----
Rearrange: A = -L^2+60L
Maximum occurs when L = -b/(2A) = -60/(2*(-1)) = 30
----
Maximum Area occurs when Length = 30 and Width = 30
=---
Maximum Area is 30^2 = 900 sq units
======================================
Cheers,
Stan H.